# No-quenching baseline for energy loss signals in oxygen-oxygen collisions

### v1

@@ Abstract
In this work, we perform computations of inclusive jet, and semi-inclusive jet-hadron cross sections for minimum bias oxygen-oxygen collisions at RHIC and LHC collision energies. We compute the no-quenching baseline for the jet nuclear modification factor $R_{\mathrm{AA}}$ and jet-, and hadron-triggered semi-inclusive nuclear modification factors $I_{\mathrm{AA}}$. We do this with state-of-the-art nuclear parton distribution functions (nPDFs), next-to-leading-order matrix elements, parton shower, and hadronization. We observe deviations from unity due to cold-nuclear matter effects, even without quenching. We demonstrate that the parton distribution uncertainties constitute a significant obstacle in detecting energy loss in small collision systems. Hadron-triggered observables are particularly sensitive to uncertainties due to correlations between the trigger and analyzed particles. For jet-triggered $I_{\mathrm{AA}}$, there exists a kinematic window in which nPDF, scale, shower, and hadronization uncertainties cancel down to percent level, overcoming a major limiting factor for energy loss discovery in small systems.
@@ 1 Introduction
A promising new direction for studying QGP in small systems is the collision of light ions, e.g., oxygen-oxygen (OO) Brewer:2021kiv or neon-neon (NeNe) Bally:2022vgo . A short run of OO collisions at $\sqrt{s_{\mathrm{NN}}}=6.8\,\text{TeV}$ at LHC is planned in 2025 Citron:2018lsq ; Bruce:2021hjk , while the STAR experiment at RHIC has already taken data at $\sqrt{s_{\mathrm{NN}}}=200\,\text{GeV}$ in 2021 Liu:2022jtl ; Huang:2023viw . Minimum bias OO collisions contain, on average, around ten participant nucleons, which is equivalent to the most peripheral lead-lead collisions Huss:2020dwe . As quenching scales with the system size Mehtar-Tani:2021fud ; Takacs:2021bpv ; Shi:2022rja , a small energy loss signal is expected in light ion collisions Huss:2020whe ; Zakharov:2021uza ; Ke:2022gkq . The successful discovery of quenching in small systems crucially depends on understanding the theory uncertainties of energy loss observables which is the main focus of this work.
@@
In this work, we study in detail the jet nuclear modification factor $R_{\rm AA}$ and the jet-, and hadron-triggered semi-inclusive nuclear modification factors $I_{\mathrm{AA}}$ in OO collisions. We recall that previous measurements of inclusive CMS:2016svx ; ATLAS:2014cpa ; ALICE:2021wct ; ALICE:2023ama and semi-inclusive ALICE:2015mdb ; ATLAS:2022iyq ; ALICE:2017svf nuclear modification factors have been used to set the upper limits on the energy loss in $p$Pb collisions. Considering both LHC and RHIC kinematics, we calculate nuclear modification factors using next-to-leading-order (NLO) matrix elements with NLO nuclear parton distributions (nPDF) and NLO-matched parton showers with hadronization. We present systematic uncertainties in all components. By neglecting any medium modification of final state interactions, we refer to this as the no-quenching baseline. Our finding shows that jet-triggered observables are much more robust than their hadronic counterpart. Furthermore, we present a kinematic window in which nPDF, scale, shower, and hadronization uncertainties cancel down to the percent level in jet-triggered $I_{\rm AA}$, overcoming a major limiting factor for energy loss discovery in small systems.
@@
The paper is organized as follows. In Section 2, we present baseline computations of jet nuclear modification factor $R_{\mathrm{AA}}$ for OO collisions at $\sqrt{s_{\mathrm{NN}}}=6.8\,\text{TeV}$ and $\sqrt{s_{\mathrm{NN}}}=200\,\text{GeV}$. We compare nPDF uncertainties from different nPDF extractions and comment on scale, shower, and hadronization uncertainties. In Section 3, we present the semi-inclusive jet-triggered nuclear modification factor and compute the baseline for the same center-of-mass energies. In particular, we analyze the (non)cancelation of nPDF uncertainties in different kinematic regions. We also study scale, parton shower, and hadronization uncertainties. In Section 4, we present the results for hadron-triggered $I_{\mathrm{AA}}$. We conclude in Section 5. A detailed description of event generator settings is included in Appendix A. Appendix B contains comparisons of computed absolute spectra for inclusive jet and jet-triggered semi-inclusive hadron cross sections with experimental data. A brief discussion of projected statistical uncertainties of the presented observables is given in Appendix C.
@@ 2 Jet nuclear modification factor $R_{\mathrm{AA}}^{j}$
We evaluate Eq. 2 with the dijet production matrix element at next-to-leading-order (NLO) using MadGraph5_aMC@NLO-v2.9.18 Alwall:2014hca combined with NLO-matched Pythia8.306 Sjostrand:2014zea ; Bierlich:2022pfr to receive showered and hadronized spectra. A detailed description of the event generator settings can be found in Appendix A. The factorization and renormalization scales were chosen to equal half the sum of the outgoing momenta $\mu_{F,R}=\frac{1}{2}\sum_{i}p_{T,i}$. After showering, we achieve NLO+leading log accuracy [^1] Pythia8 is almost next-to-leading-log accurate, see Ref. Dasgupta:2020fwr for details.. We will use nPDF parametrizations implemented in the LHAPDF6 Buckley:2014ana framework. Jets were clustered on all final state particles, using the anti-$k_{T}$ algorithm in fastjet Cacciari:2008gp ; Cacciari:2011ma with $R=0.4$ and $|y_{\mathrm{jet}}|<3$. At NLO accuracy, the inclusive jet spectra reach excellent agreement with experimental data from $pp$ collision, see in Appendix B.
@@
Figure 1: Predictions for the no-quenching baseline for jet nuclear modification factor in minimum bias oxygen-oxygen collisions at $\sqrt{s_{\mathrm{NN}}}=6.8\,\text{TeV}$. Solid bands show 9-point scale variation and hatched bands indicate 90% confidence interval of nNPDF uncertainties. Left: Comparison of EPPS16 and EPPS21. Right: Comparison of EPPS21 and TUJU21.
The left panel of Fig. 1 shows $R_{\mathrm{AA}}^{j}$ as a function of jet $p_{T}$ and compares results obtained with EPPS21 Eskola:2021nhw to the older EPPS16 Eskola:2016oht nPDF set. Correspondingly the $pp$ reference cross section is calculated with CT18ANLO Hou:2019efy and CT14NLO Dulat:2015mca free proton PDFs. The range of experimental data to constrain nPDFs in EPPS21 was extended to also contain recent DIS data, W±, dijet and D0 production. No final state interactions are included in our calculation, nevertheless the central value of $R_{\rm AA}\neq 1$ because of the nPDFs. The nPDF uncertainty of the two parametrizations differ especially at low $p_{T}$, where EPPS21 is four times more uncertain. Larger uncertainties could originate from the updated $A$-dependence of the parametrization. The functional form in EPPS21 includes additional parameters (24 instead of 20), making the fit more flexible. The measurement of dijet production in proton-oxygen collisions could be used to further constrain the nPDF uncertainties for light ions Paakkinen:2021jjp . Uncertainties coming from scale variations of $\mu_{F}$ and $\mu_{R}$ are negligible compared to the nPDF uncertainties.
The right panel of Fig. 1 compares EPPS21 and TUJU21 Helenius:2021tof nPDFs. While central values are quite different, TUJU21 predictions are mostly contained within the large EPPS21 uncertainty band [^2] The sudden increase in uncertainties at 45-50 GeV bin for TUJU21 is a numerical effect and should be disregarded.. We note that TUJU21 extractions use only 16 free parameters in contrast to 24 of EPPS21. Therefore, smaller uncertainties could result from more restrictive parametrization.
@@
Baseline calculations for charged jet $R_{\mathrm{AA}}^{j}$ at $\sqrt{s_{\mathrm{NN}}}=200\,\text{GeV}$ are shown in Fig. 3. In this case, the clustering was performed on charged particles, requiring $|\eta_{\mathrm{jet}}|<0.6$. Deviation of $R_{\mathrm{AA}}^{j}$ from unity is clearly present due to the nPDFs. As in Fig. 1, TUJU21 tends to have smaller uncertainties than EPPS21, except for large momentum. The central values differ among nPDF sets but they still agree within the errorbands. For discussion of the seen negative trend in $R_{\mathrm{AA}}^{j}$ see Section 3.3.
@@ 3.1 No-quenching baseline results
The four panels in Fig. 4 compare the no-quenching baseline of $I_{\mathrm{OO}}^{h}$ for near- and away-side yields obtained with different oxygen nPDFs and for different jet-momentum triggers at 6.8 TeV center-of-mass energy [^3] When comparing computed total yields $Y_{\mathrm{AA}}^{h}$ with measurements in $pp$ collisions in Appendix B, we find that the computations over predict the amount of produced hadrons but this difference in magnitude cancels after taking the ratio to construct the semi-inclusive modification factor $I_{\mathrm{AA}}^{h}$.. The central values deviate from unity because of the nPDFs and they vary for different nPDF parametrizations at high $p_{T}$. EPPS21 results tend to become larger, while TUJU21 is moderately smaller than unity. The nPDF uncertainties are smaller than those for jet $R_{\rm AA}$ in Fig. 1. In some kinematic region, they are ten times smaller, demonstrating the resilience of the double ratio structure. As $p_{T}$ becomes larger, nPDF uncertainties show significant growth. The characteristics of near- and away-side $I_{\mathrm{OO}}^{h}$ are similar (upper versus lower panels). Higher trigger jet momenta (right panels), significantly decrease nPDF uncertainties for shown $p_{T}$. In Section 3.2 we give a detailed explanation for the trends seen in uncertainty cancelation and baseline deviation from unity.
@@
3.2 Cancelation of nPDF uncertainty and deviation from unity
@@
To gain a simple intuition in uncertainty cancelation, consider a more straightforward case where uncertainties can be computed from a variance over the members of the error set, i.e., $\Delta X\equiv\mathrm{var}(X[f])$. Then the uncertainty of the ratio $X/Y$ is given by
@@
We see that the cancelation of uncertainties in the ratio $X/Y$ is only possible when the correlation is positive $\rho>0$. If $\rho<0$, we have the case when the uncertainties in the ratio even increase.
Because $Y_{pp}$ does not depend on the nPDF, the uncertainty of $I_{\rm AA}$ is especially sensitive to the cancelation of nPDF uncertainties between $\sigma^{j}_{\rm AA}$ and $\sigma^{h+j}_{\rm AA}$. To study this, we compute the Pearson correlation,
@@
using the members of nPDF uncertainty sets [^4] Variance of Hessian PDF error sets does not give the correct uncertainty, cf. Eq. 8, therefore our results are only indicative. In addition, nPDF error sets also include uncertainties due to the proton reference, but this uncertainty cancels in the ratio between $Y_{\rm AA}$ and $Y_{pp}$ to a large degree.. Figure 6 shows the Pearson correlation for LHC kinematics (RHIC kinematics would conclude the same). The nPDF uncertainties are strongly correlated below 30 GeV, explaining the observed uncertainty cancelation in Figs. 4 and 5. As $p_{T}$ approaches $p_{T,\mathrm{min}}^{j}$, the correlation decreases visibly increasing the nPDF uncertainty. Increasing $p^{j}_{T,\mathrm{min}}$ on the right panel of Fig. 6, we find large positive correlation over a wider range of hadronic transverse momenta, consistent with broad uncertainty cancelation visible in the right panels of Fig. 4.
@@
In the following, we will use parton kinematics to explain the loss of uncertainty correlations between high momentum hadrons and trigger jet cross section. The nPDFs in factorization formula Eq. 2 are evaluated at Bjorken $x_{1,2}$, where at leading order we have Ellis:318585
@@
Here, $p_{T}$ and $y_{1,2}$ denote the transverse momentum and rapidity of outgoing back-to-back partons. Figure 7 shows the distribution of $x_{1,2}$ values for different event selections relevant for $I_{\rm AA}$. We used Pythia8 with EPPS21 NLO nPDFs (as LO ones are not available). The left panel shows the probed $x_{1,2}$ values in the inclusive jet selection for $p_{T}^{j}>30\,\text{GeV}$. The steeply falling jet spectrum biases the $x$ values towards the bound controlled by $p_{T,\text{min}}$, Eq. 12 (dotted line). The middle panel shows $h+j$ events for low momentum hadrons $p_{T}<10\,\text{GeV}$ that comes from fragmentation of $p_{T}\approx p_{T,\text{min}}$ jets. Therefore, the probed $x_{1,2}$ in these selections are similar to those probed by inclusive jets and nPDF uncertainties are positively correlated in Fig. 6. The rightmost panel is for high momentum selection (hadrons with $p_{T}>40\,\text{GeV}$). This time hadrons must come from a larger momentum jet with $p_{T}>p_{T,\text{min}}$. Correspondingly, a high momentum hadron cross section corresponds to larger values of Bjorken-$x$ and is less correlated with an inclusive jet cross section as seen in Fig. 6.
The difference in the probed Bjorken-$x$ by the trigger and semi-inclusive cross sections can be also used to explain the observed $I_{\rm AA}$ deviation from unity. Depending on the values of Bjorken-$x$, parton distributions inside the nuclei (and cross sections) are suppressed/enhanced compared with a free proton. This can be seen in the inclusive jet $R_{\rm AA}$, cf., Figs. 1 and 3. For very small momentum (Bjorken-$x$), we expect the gluon shadowing to suppress the jet production, while for the intermediate momentum, it is enhanced by anti-shadowing. For large jet momentum (Bjorken-$x$), the parton distributions and, hence, cross sections are suppressed by the EMC effect. Such a rising and then falling trend is well seen for EPPS21 results in Fig. 1. For TUJU21, we observe a mostly falling trend. Note that at lower collision energies shown in Fig. 3, the probed Bjorken-$x$ is larger and we observe the decreasing $R_{\rm AA}$ with momentum. The trigger jet cross section is dominated by jets of momentum just above the threshold $p_{T,\text{min}}^{j}$, therefore we can approximate $\left.\sigma^{j}\right|_{p_{T}^{j}>p_{T,\text{min}}^{j}}=\int_{p_{T,\text{min}}^{j}}^{\infty}\mathrm{d}p_{T}\,\mathrm{d}\sigma^{j}(p_{T})/\mathrm{d}p_{T}\approx\Delta p^{j}_{T,\text{min}}\mathrm{d}\sigma^{j}(p^{j}_{T,\text{min}})/\mathrm{d}p_{T}$. Using this, the semi-inclusive nuclear modification factor can be written as the ratio of two nuclear modification factors
@@
As shown above, $h+j$ cross section with hadron momentum above $p_{T,\text{min}}^{j}$ probes larger values of Bjorken-$x$ than the trigger. For EPPS21 nPDFs, jet $R^{j}_{\mathrm{AA}}$ is increasing in the region $40\,\text{GeV}<p_{T}<200\,\text{GeV}$, while for TUJU21 it is constant or decreasing. We expect that hadrons with momentum $p_{T}<90\,\text{GeV}$ will be produced mostly by jets with $p_{T}<200\,\text{GeV}$. Then $I_{\mathrm{AA}}$ will be larger or smaller than unity, depending if $R^{j}_{\mathrm{AA}}$ increases or decreases above the trigger momentum. Indeed in Fig. 4 we see that for large hadron momentum $I_{\mathrm{AA}}>1$ for EPPS21 and $I_{\mathrm{AA}}<1$ for TUJU21 nPDFs, in agreement with our qualitative argument. $R^{j}_{\mathrm{AA}}$ has negative slope at $\sqrt{s_{\rm NN}}=200\,\text{GeV}$ shown in Fig. 3 for both nPDF sets and indeed we see strongly suppressed $I_{\mathrm{AA}}<1$ in Fig. 5.

### v3

@@ Abstract
In this work, we perform computations of inclusive jet, and semi-inclusive jet-hadron cross sections for minimum bias oxygen-oxygen collisions at RHIC and LHC collision energies. We compute the no-quenching baseline for the jet nuclear modification factor $R_{\mathrm{AA}}$ and jet-, and hadron-triggered semi-inclusive nuclear modification factors $I_{\mathrm{AA}}$. We do this with state-of-the-art nuclear parton distribution functions (nPDFs), next-to-leading-order matrix elements, parton shower, and hadronization. We observe deviations from unity due to cold-nuclear matter effects, even without quenching. We demonstrate that the parton distribution uncertainties constitute a significant obstacle in detecting energy loss in small collision systems. Hadron-triggered observables are particularly sensitive to uncertainties due to correlations between the trigger and analyzed particles. For jet-triggered $I_{\mathrm{AA}}$, there exists a kinematic window in which nPDF and scale uncertainties cancel dramatically while showing little sensitivity to parton shower and hadronization models, addressing a major limiting factor for energy loss discovery in small systems.
@@ 1 Introduction
A promising new direction for studying QGP in small systems is the collision of light ions, e.g., oxygen-oxygen (OO) Brewer:2021kiv or neon-neon (NeNe) Bally:2022vgo . A short run of OO collisions at $\sqrt{s_{\mathrm{NN}}}=6.8\,\text{TeV}$ at LHC is planned in 2025 Citron:2018lsq ; Bruce:2021hjk , while the STAR experiment at RHIC has already taken data at $\sqrt{s_{\mathrm{NN}}}=200\,\text{GeV}$ in 2021 Liu:2022jtl ; Huang:2023viw . Minimum bias OO collisions contain, on average, around ten participant nucleons, which is equivalent to the most peripheral lead-lead collisions Huss:2020dwe . As quenching scales with the system size Mehtar-Tani:2021fud ; Takacs:2021bpv ; Mehtar-Tani:2024jtd ; Shi:2022rja , a small energy loss signal is expected in light ion collisions Katz:2019qwv ; Huss:2020whe ; Zakharov:2021uza ; Xie:2022fak ; Ke:2022gkq ; Vitev:2023nti ; Ogrodnik:2024qug . The successful discovery of quenching in small systems crucially depends on understanding the theory uncertainties of energy loss observables which is the main focus of this work.
@@
In this work, we study in detail the jet nuclear modification factor $R_{\rm AA}$ and the jet-, and hadron-triggered semi-inclusive nuclear modification factors $I_{\mathrm{AA}}$ in OO collisions. We recall that previous measurements of inclusive CMS:2016svx ; ATLAS:2014cpa ; ALICE:2021wct ; ALICE:2023ama and semi-inclusive ALICE:2015mdb ; ATLAS:2022iyq ; ALICE:2017svf nuclear modification factors have been used to set the upper limits on the energy loss in $p$Pb collisions. Considering both LHC and RHIC kinematics, we calculate nuclear modification factors using next-to-leading-order (NLO) matrix elements with NLO nuclear parton distributions (nPDF) and NLO-matched parton showers with hadronization. We present systematic uncertainties in nPDF and scale components as well as sensitivity to parton showers and hadronization models. By neglecting any medium modification of final state interactions, we refer to this as the no-quenching baseline. Our finding shows that jet-triggered observables are much more robust than their hadronic counterpart. Furthermore, we present a kinematic window in which nPDF and scale uncertainties cancel dramatically in jet-triggered $I_{\rm AA}$ while showing little sensitivity to parton shower and hadronization models, addressing a major limiting factor for energy loss discovery in small systems.
@@
The paper is organized as follows. In Section 2, we present baseline computations of jet nuclear modification factor $R_{\mathrm{AA}}$ for OO collisions at $\sqrt{s_{\mathrm{NN}}}=6.8\,\text{TeV}$ and $\sqrt{s_{\mathrm{NN}}}=200\,\text{GeV}$. We compare nPDF uncertainties from different nPDF extractions and comment on scale, shower, and hadronization uncertainties. In Section 3, we present the semi-inclusive jet-triggered nuclear modification factor and compute the baseline for the same center-of-mass energies. In particular, we analyze the (non)cancellation of nPDF uncertainties in different kinematic regions. We also study scale, parton shower, and hadronization uncertainties. In Section 4, we present the results for hadron-triggered $I_{\mathrm{AA}}$. We conclude in Section 5. A detailed description of event generator settings is included in Appendix A. Appendix C contains comparisons of computed absolute spectra for inclusive jet and jet-triggered semi-inclusive hadron cross sections with experimental data. A brief discussion of projected statistical uncertainties of the presented observables is given in Appendix D.
@@ 2 Jet nuclear modification factor $R_{\mathrm{AA}}^{j}$
We evaluate Eq. 2 with the dijet production matrix element at next-to-leading-order (NLO) using MadGraph5_aMC@NLO-v2.9.18 Alwall:2014hca combined with NLO-matched Pythia8.306 Sjostrand:2014zea ; Bierlich:2022pfr to receive showered and hadronized spectra. A detailed description of the event generator settings can be found in Appendix A. The factorization and renormalization scales were chosen to equal half the sum of the outgoing momenta $\mu_{F,R}=\frac{1}{2}\sum_{i}p_{T,i}$. After showering, we achieve NLO+leading log accuracy. We will use nPDF parametrizations implemented in the LHAPDF6 Buckley:2014ana framework. Jets were clustered on all final state particles, using the anti-$k_{T}$ algorithm in fastjet Cacciari:2008gp ; Cacciari:2011ma with $R=0.4$ and $|y_{\mathrm{jet}}|<3$. At NLO accuracy, the inclusive jet spectra reach excellent agreement with experimental data from $pp$ collision, see in Appendix C.
@@
Figure 1: Predictions for the no-quenching baseline for jet nuclear modification factor in minimum bias oxygen-oxygen collisions at $\sqrt{s_{\mathrm{NN}}}=6.8\,\text{TeV}$. Solid bands show 9-point scale variation and hatched bands indicate 90% confidence interval of nPDF uncertainties. Left: Comparison of EPPS16 and EPPS21. Right: Comparison of EPPS21, TUJU21 and nNNPDF3.0.
The left panel of Fig. 1 shows $R_{\mathrm{OO}}^{j}$ as a function of jet $p_{T}$ and compares results obtained with EPPS21 Eskola:2021nhw to the older EPPS16 Eskola:2016oht nPDF set. Correspondingly the $pp$ reference cross section is calculated with CT18ANLO Hou:2019efy and CT14NLO Dulat:2015mca free proton PDFs. The range of experimental data to constrain nPDFs in EPPS21 was extended to also contain recent DIS data, W±, dijet and D0 production. No final state interactions are included in our calculation, nevertheless the central value of $R_{\rm OO}\neq 1$ because of the nPDFs. The nPDF uncertainty of the two parametrizations differ especially at low $p_{T}$, where EPPS21 is four times more uncertain. Larger uncertainties could originate from the updated $A$-dependence of the parametrization. The functional form in EPPS21 includes additional parameters (24 instead of 20), making the fit more flexible. The measurement of dijet production in proton-oxygen collisions could be used to further constrain the nPDF uncertainties for light ions Paakkinen:2021jjp . Uncertainties coming from scale variations of $\mu_{F}$ and $\mu_{R}$ are negligible compared to the nPDF uncertainties.
The right panel of Fig. 1 compares EPPS21, TUJU21 Helenius:2021tof and nNNPDF3.0 AbdulKhalek:2022fyi nPDFs. We observed occasional numerical instabilities in evaluating NLO uncertainties, leading to a sudden increase, as exemplified in the 45-50 GeV bin for TUJU21. Consequently, for nNNPDF3.0 we only show results obtained with matrix elements at leading-order (LO) combined with the NLO nPDF. Comparisons between LO and NLO matrix elements combined with the NLO nPDF results in almost identical $R_{\rm AA}$ (and $I_{\rm AA}$) nPDF uncertainty bands, see Appendix B. While central values are quite different, TUJU21 and nNNPDF3.0 predictions are mostly contained within the EPPS21 uncertainty band below $p_{T}=200\,\mathrm{GeV}$. Above, TUJU21 and nNNPDF3.0 uncertainties become larger but all predictions are still consistent with each other. We note that TUJU21 extractions use only 16 free parameters in contrast to 24 of EPPS21. Therefore, smaller uncertainties could result from more restrictive parametrization. On the other hand, nNNPDF3.0 uses a neural network-based approach that can incorporate a much larger number of parameters and fewer assumptions on the functional form of the nPDFs. Therefore nNNPDF3.0 uncertainties can grow rapidly in kinematic regions not constrained by experimental data.
@@
Baseline calculations for charged jet $R_{\mathrm{OO}}^{j}$ at $\sqrt{s_{\mathrm{NN}}}=200\,\text{GeV}$ are shown in Fig. 3. In this case, the clustering was performed on charged particles, requiring $|\eta_{\mathrm{jet}}|<0.6$. Deviation of $R_{\mathrm{OO}}^{j}$ from unity is clearly present due to the nPDFs. As in Fig. 1, TUJU21 tends to have smaller uncertainties than EPPS21, except for large momentum. nNNPDF3.0 errors are large across all shown $p_{T}$. The central values differ among nPDF sets but still agree within errorbands. For discussion of the seen negative trend in $R_{\mathrm{OO}}^{j}$ see Section 3.3.
@@ 3.1 No-quenching baseline results
The four panels in Fig. 4 compare the no-quenching baseline of $I_{\mathrm{OO}}^{h}$ for near- and away-side yields obtained with different oxygen nPDFs and for different jet-momentum triggers at 6.8 TeV center-of-mass energy [^1] When comparing computed total yields $Y_{\mathrm{AA}}^{h}$ with measurements in $pp$ collisions in Appendix C, we find that the computations over predict the amount of produced hadrons but this difference in magnitude cancels after taking the ratio to construct the semi-inclusive modification factor $I_{\mathrm{AA}}^{h}$.. The central values deviate from unity because of the nPDFs and they vary for different nPDF parametrizations at high $p_{T}$. EPPS21 results tend to become larger, while TUJU21 and nNNPDF3.0 are moderately smaller than unity. The nPDF uncertainties are smaller than those for jet $R_{\rm OO}$ in Fig. 1. In some kinematic region, they are ten times smaller, demonstrating the resilience of the double ratio structure. As $p_{T}$ becomes larger, nPDF uncertainties show significant growth. The characteristics of near- and away-side $I_{\mathrm{OO}}^{h}$ are similar (upper versus lower panels). Higher trigger jet momenta (right panels), significantly decrease nPDF uncertainties for shown $p_{T}$. In Section 3.2 we give a detailed explanation for the trends seen in uncertainty cancellation and baseline deviation from unity.
@@
3.2 Cancellation of nPDF uncertainty and deviation from unity
@@
The neural network-based nNNPDF framework uses the replica method Forte:2002fg . The error set corresponds to the Monte Carlo sample of the probability density of the fitted data. The $X\%$ confidence interval of some observable $X[f]$ is computed by sorting the values of the error set $X[f_{i}]$ and removing symmetrically $(100-X)\%$ members of the replicas with the highest and lowest values. The central value is the median value of the error set.
To gain a simple intuition in uncertainty cancellation, consider a more straightforward case where uncertainties can be computed from a variance over the members of the error set, i.e., $\Delta X\equiv\mathrm{var}(X[f])$. Then the uncertainty of the ratio $X/Y$ is given by
@@
We see that the cancellation of uncertainties in the ratio $X/Y$ is only possible when the correlation is positive $\rho>0$. If $\rho<0$, we have the case when the uncertainties in the ratio even increase.
Because $Y_{pp}$ does not depend on the nPDF, the uncertainty of $I_{\rm AA}$ is especially sensitive to the cancellation of nPDF uncertainties between $\sigma^{j}_{\rm AA}$ and $\sigma^{h+j}_{\rm AA}$. To study this, we compute the Pearson correlation,
@@
using the members of nPDF uncertainty sets [^2] Variance of Hessian PDF error sets does not give the correct uncertainty, cf. Eq. 8, therefore our results are only indicative. In addition, nPDF error sets also include uncertainties due to the proton reference, but this uncertainty cancels in the ratio between $Y_{\rm AA}$ and $Y_{pp}$ to a large degree.. Figure 6 shows the Pearson correlation for LHC kinematics (RHIC kinematics would conclude the same). The nPDF uncertainties are strongly correlated below 30 GeV, explaining the observed uncertainty cancellation in Figs. 4 and 5. As $p_{T}$ approaches $p_{T,\mathrm{min}}^{j}$, the correlation decreases visibly increasing the nPDF uncertainty. Increasing $p^{j}_{T,\mathrm{min}}$ on the right panel of Fig. 6, we find large positive correlation over a wider range of hadronic transverse momenta, consistent with broad uncertainty cancellation visible in the right panels of Fig. 4.
@@
In the following, we will use parton kinematics to explain the loss of uncertainty correlations between high momentum hadrons and trigger jet cross section. The nPDFs in the factorization formula Eq. 2 are evaluated at Bjorken $x_{1,2}$, where at leading order we have Ellis:318585
@@
Here, $p_{T}$ and $y_{1,2}$ denote the transverse momentum and rapidity of outgoing back-to-back partons. Figure 7 shows the distribution of $x_{1,2}$ values for different event selections relevant for $I_{\rm AA}$. We used Pythia8 with LO matrix element and EPPS21 NLO nPDFs (as LO ones are not available). We used the Pythia8::Info::x1pdf() function to access $x_{1,2}$ at the highest scale. We neglect the $x_{1,2}$ sampling during initial state radiation for simplification and better agreement with the strict LO picture. The left panel shows the probed $x_{1,2}$ values in the inclusive jet selection for $p_{T}^{j}>30\,\text{GeV}$. The steeply falling jet spectrum biases the $x$ values towards the bound controlled by $p_{T,\text{min}}$, Eq. 12 (dotted line). The middle panel shows $h+j$ events for low momentum hadrons $p_{T}<10\,\text{GeV}$ that comes from fragmentation of $p_{T}\approx p_{T,\text{min}}$ jets. Therefore, the probed $x_{1,2}$ in these selections are similar to those probed by inclusive jets and nPDF uncertainties are positively correlated in Fig. 6. The rightmost panel is for high momentum selection (hadrons with $p_{T}>40\,\text{GeV}$). This time hadrons must come from a larger momentum jet with $p_{T}>p_{T,\text{min}}$. Correspondingly, a high momentum hadron cross section corresponds to larger values of Bjorken-$x$ and is less correlated with an inclusive jet cross section as seen in Fig. 6.
The difference in the probed Bjorken-$x$ by the trigger and semi-inclusive cross sections can be also used to explain the observed $I_{\rm OO}$ deviation from unity. Depending on the values of Bjorken-$x$, parton distributions inside the nuclei (and cross sections) are suppressed/enhanced compared with a free proton. This can be seen in the inclusive jet $R_{\rm OO}$, cf., Figs. 1 and 3. For very small momentum (Bjorken-$x$), we expect the gluon shadowing to suppress the jet production, while for the intermediate momentum, it is enhanced by anti-shadowing. For large jet momentum (Bjorken-$x$), the parton distributions and, hence, cross sections are suppressed by the EMC effect. Such a rising and then falling trend is well seen for EPPS21 results in Fig. 1. For TUJU21 and nNNPDF3.0, we observe a mostly falling trend. Note that at lower collision energies shown in Fig. 3, the probed Bjorken-$x$ is larger and we observe the decreasing $R_{\rm OO}$ with momentum. The trigger jet cross section is dominated by jets of momentum just above the threshold $p_{T,\text{min}}^{j}$, therefore we can approximate $\left.\sigma^{j}\right|_{p_{T}^{j}>p_{T,\text{min}}^{j}}=\int_{p_{T,\text{min}}^{j}}^{\infty}\mathrm{d}p_{T}\,\mathrm{d}\sigma^{j}(p_{T})/\mathrm{d}p_{T}\approx\Delta p^{j}_{T,\text{min}}\mathrm{d}\sigma^{j}(p^{j}_{T,\text{min}})/\mathrm{d}p_{T}$. Using this, the semi-inclusive nuclear modification factor can be written as the ratio of two nuclear modification factors
@@
As shown above, $h+j$ cross section with hadron momentum above $p_{T,\text{min}}^{j}$ probes larger values of Bjorken-$x$ than the trigger. For EPPS21 nPDFs, jet $R^{j}_{\mathrm{OO}}$ is increasing in the region $40\,\text{GeV}<p_{T}<200\,\text{GeV}$, while for TUJU21 it is constant or decreasing. We expect that hadrons with momentum $p_{T}<90\,\text{GeV}$ will be produced mostly by jets with $p_{T}<200\,\text{GeV}$. Then $I_{\mathrm{OO}}$ will be larger or smaller than unity, depending if $R^{j}_{\mathrm{OO}}$ increases or decreases above the trigger momentum. Indeed in Fig. 4 we see that for large hadron momentum $I_{\mathrm{OO}}>1$ for EPPS21 and $I_{\mathrm{OO}}<1$ for TUJU21 nPDFs, in agreement with our qualitative argument. nNNPDF3.0 jet nuclear modification factors remain centered around unity which results in relatively constant central $I_{\mathrm{OO}}^{h}$. $R^{j}_{\mathrm{OO}}$ has negative slope at $\sqrt{s_{\rm NN}}=200\,\text{GeV}$ shown in Fig. 3 for both nPDF sets and indeed we see strongly suppressed $I_{\mathrm{OO}}<1$ in Fig. 5.